OpenStudy (anonymous):

a lab network consisting of 20 computers was attack by a computer virus. this virus enters each computer with probability 0.4 independent of others. find the probability that it enters atleast 10 computers? its answer is 0.2447 but i dont know which distribution i should use ?

OpenStudy (cwrw238):

i think you can use the Binomial Distribution

OpenStudy (anonymous):

but i gives 0.1171

OpenStudy (cwrw238):

oh - i'm pretty rusty at this stuff - did you use Binomial tables?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

ok let me see the table

OpenStudy (cwrw238):

i'll see if i can dig mine out

OpenStudy (anonymous):

its of no use

OpenStudy (anonymous):

the table is giving .8725

OpenStudy (cwrw238):

P(X>=10) = 1 -P(X<=9) = 1 - F(9;20,0.4)

OpenStudy (cwrw238):

i cant find my tables i'm afraid

OpenStudy (anonymous):

np

OpenStudy (anonymous):

@cwrw238 can you give the half solution i m again confused plz

OpenStudy (cwrw238):

ah - the tables are online P = 1 - 0.755 = 0.245 according to them

OpenStudy (anonymous):

actually the thing is that i m noob to probability

OpenStudy (cwrw238):

i used the cumulative binomial table

OpenStudy (anonymous):

so thanks alot for your help

OpenStudy (cwrw238):

yw - i'm not great at probability - don;t like it very much

OpenStudy (anonymous):

@cwrw238 tell me how i can improve my probability

OpenStudy (cwrw238):

i'm sorry i cant really help you there - i expect one of the legends, senseis or gurus should be able to help

OpenStudy (anonymous):

You can use the Standard Model Equation...